an=-an-1+16an-2+4an-3-48an-4,where a°=0,a1=16,a2=-2,a3=142
1
Expert's answer
2021-04-13T13:58:07-0400

Let us solve "a_n=-a_{n-1}+16a_{n-2}+4a_{n-3}-48a_{n-4}," where "a_0=0,a_1=16,a_2=-2,a_3=142".


Firstly, let us solve the characteristic equation


"k^4=-k^3+16k^2+4k-48"


"k^4+k^3-16k^2-4k+48=0"


"k^4-4k^2+k^3-4k-12k^2+48=0"


"k^2(k^2-4)+k(k^2-4)-12(k^2-4)=0"


"(k^2+k-12)(k^2-4)=0"


"(k-3)(k+4)(k-2)(k+2)=0"


It follows that the characteristoc equation has the following roots:


"k_1=3,\ k_2=-4,\ k_3=2,\ k_4=-2."


Therefore, the general solution is "a_n=c_13^n+c_2(-4)^n+c_32^n+c_4(-2)^n."


Since "a_0=0,a_1=16,a_2=-2,a_3=142", we conclude that


"\begin{cases}
0=a_0=c_1+c_2+c_3+c_4\\
16=a_1=3c_1-4c_2+2c_3-2c_4\\
-2=a_2=9c_1+16c_2+4c_3+4c_4\\
142=a_3=27c_1-64c_2+8c_3-8c_4
\end{cases}"


"\begin{cases}
c_1+c_2+c_3+c_4=0\\
3c_1-4c_2+2c_3-2c_4=16\\
5c_1+12c_2=-2\\
15c_1-48c_2=78
\end{cases}"


"\begin{cases}
1+c_3+c_4=0\\
10+2c_3-2c_4=16\\
c_2=-1\\
c_1=2
\end{cases}"


"\begin{cases}
c_4=-2\\
c_3=1\\
c_2=-1\\
c_1=2
\end{cases}"


It follows that the final solusion is


"a_n=2\cdot3^n-(-4)^n+2^n-2\cdot (-2)^n."